Determine the maximum voltage drop between a and c points in the circuit. d 1uF Ο Ο 55V Ο Ο 38.8V Ο Ο 116.5V Ο Ο 86.8V
Answer #1

IK lui. OTO- v=140 sin (5oot) 5002 IHSG LUE Xi = WL = 5008 Iz Xc 2 | 3 | - wc Joox 10-6 R = 1000 - 20002 Z-R2Cxe-x2)2 = (1000IK lui. OTO- v=140 sin (5oot) 5002 IHSG LUE Xi = WL = 5008 Iz Xc 2 | 3 | - wc Joox 10-6 R = 1000 - 20002 Z-R2Cxe-x2)2 = (1000